OpenStudy (anonymous):

What is the side length of a cube with a volume of 2744 cm3?

OpenStudy (anonymous):

V = 2744 mm^3 Let side of cube = a V = a^3 => 2744 mm^3 = a^3 => (14 mm)^3 = a^3 => a = 14 mm

OpenStudy (anonymous):

So the answer is 14?

OpenStudy (anonymous):

Wait no that can't be right..

OpenStudy (anonymous):